How do you find the coordinates of the vertex?
heres a problem y=3x^2

Answer by stanbon(75887)

Rewrite the equation in the form y-k=3(x-h)^2
The (h,k) will be the vertex.
Your problem becomes y-0=3(x-0)^2
So, h=0 and k=0.
Cheers,
Stan H.